OpenStudy (kaylala):

Trigonometric Equation tanA cosA + √(3) cosA = 0

OpenStudy (anonymous):

\[ \tan(A) \cos(A) + \sqrt 3 \cos(A)=0\\ \cos(A) (\tan(A) ) + \sqrt 3 )=0 \] Can you finish it now?

OpenStudy (anonymous):

NOw you have two equations to solve \[ \cos(A) =0 \\ \text { or }\\ \tan(A) + \sqrt 3 =0 \]

OpenStudy (anonymous):

So \[ A= \frac \pi 2 + k \, \pi\\ A= -\frac \pi 3 + k \,\pi \]

OpenStudy (kaylala):

please simplify. i'm new to this lesson i really dont get you

OpenStudy (anonymous):

Do you know when cos(x)=0?

OpenStudy (kaylala):

nope

OpenStudy (anonymous):

Try to go back to your book and study the function cos and tan etc before.

OpenStudy (kaylala):

i dont even know how to start solving this